# Machine learning many-body potentials for charged colloids in primitive 1:1 electrolytes

### v1

@@ Abstract
Effective interactions between charged particles dispersed in an electrolyte are most commonly modeled using the Derjaguin-Landau-Verwey-Overbeek (DLVO) potential, where the ions in the suspension are coarse-grained out at mean-field level. However, several experiments point to shortcomings of this theory, as the distribution of ions surrounding colloids is governed by nontrivial correlations in regimes of strong Coulomb coupling (e.g. low temperature, low dielectric constant, high ion valency, high surface charge). Insight can be gained by explicitly including the ions in simulations of these colloidal suspensions, even though direct simulations of dispersions of highly charged spheres are computationally demanding. To circumvent slow equilibration, we employ a machine-learning (ML) framework to generate ML potentials that accurately describe the effective colloid interactions. These ML potentials enable fast simulations and make large-scale simulations of charged colloids in suspension possible, opening the possibility for a systematic study of their phase behaviour, in particular gas-liquid and fluid-solid coexistence.
@@ I Introduction
When the refractive index of the solvent is matched to that of the colloidal particles, the attractive part of the DLVO potential, arising from Van der Waals forces due to dipole-dipole fluctuations, can be neglected. Under these conditions, the interaction between two particles with the same charge sign reduces to a purely repulsive electrostatic potential. It was therefore surprising when a series of experiments reported attractive interactions between like-charged colloids in systems of highly charged, micron-scale particles at low-salt concentrations,Kepler and Fraden (1994); Ito, Yoshida, and Ise (1994); Tata et al. (1997); Larsen and Grier (1997); Gomez et al. (2009); Wang et al. (2024) challenging the universal validity of DLVO theory. It should be noted, however, that the results of some of these experiments remain controversial, due to issues related to reproducibility Palberg and Würth (1994) and potential experimental artifacts.Squires and Brenner (2000) This low-salt regime is characterized by long screening lengths, which increases the importance of many-body interactions beyond simple pairwise interactions, as the larger electric double layers more frequently overlap with multiple neighboring EDLs. A second characteristic of this regime is that, as the colloid concentration increases, the ionic strength, and thus the effective Debye length, is no longer determined predominantly by the concentration of the background electrolyte, but rather by the higher concentration of counterions. Various studies based on Poisson-Boltzmann calculations have proposed that these two effects are the leading causes behind like-charge attractions—specifically, gas-liquid and gas-crystal phase separation—in low-salt colloidal suspensions.van Roij and Hansen (1997); van Roij, Dijkstra, and Hansen (1999); Zoetekouw and van Roij (2006); Warren (2000); Russ et al. (2002) A second cause of like-charged attractive interactions lies in the strongly coupled electrostatic regime, which is characterized by low temperature, low dielectric constant, and high ionic valency.Kjellander, Marčelja, and Quirk (1988); Allahyarov, D’Amico, and Löwen (1998); Naji et al. (2005); Punkkinen et al. (2008) A defining feature of this regime is the binding of point-like ions onto highly charged surfaces, forming a quasi-two-dimensional layer rather than the diffuse three-dimensional distribution characteristic of the Poisson-Boltzmann electric double layer. When two such decorated surfaces approach each other, strong ion-ion correlations give rise to attractive interactions.
@@
In this paper, we also coarse-grain the ions to significantly reduce simulation times, but we achieve this by employing a machine learning (ML) framework to generate effective many-body colloidal potentials. While ML approaches have been successfully used to accelerate simulations of uncharged systems,Deringer, Caro, and Csányi (2019); Campos-Villalobos et al. (2021); Nguyen and Huang (2022); Argun, Fu, and Statt (2024) the present work is the first to apply machine-learned effective potentials to charged colloids. To construct these ML colloidal potentials, we adopt a linear regression scheme previously used to develop effective potentials for sterically stabilized colloids with non-adsorbing polymers Campos-Villalobos et al. (2021) and for ligand-stabilized colloids.Giunta, Campos-Villalobos, and Dijkstra (2023) Here, we train the ML potentials using data from primitive model simulations, resulting in effective potentials that accurately capture the effective interactions between charged colloids in the presence of ions, in addition to bare-colloid interactions. In subsequent “colloids-only” simulations using these ML potentials, the ions are effectively integrated out. Moreover, in this ML potential representation, the colloids are essentially charge-neutral, thus eliminating the need for considering long-range interactions. These two factors bring about a significant increase in computational efficiency –at least twenty times faster for colloidal systems–, enabling studies at higher colloid charges and salt concentrations.
@@ II.1 Primitive Model simulations
The system of interest in this study is a three-dimensional primitive model (PM) of a colloidal dispersion in a 1:1 electrolyte. This colloidal suspension is viewed as a three-component mixture of charged colloids, monovalent counterions, and monovalent coions, while the solvent is modeled as a structureless dielectric continuum. The colloidal particles have an effective hard-core diameter $\sigma_{c}$ and a total positive charge $Q_{c}q$ uniformly distributed over their surfaces, with valency $Q_{c}\geq 1$ and elementary charge $q$. The monovalent counterions are anions with diameter $\sigma_{-}$ and charge $-q$, and the monovalent coions are cations of diameter $\sigma_{+}$ and charge $+q$. These three species are illustrated in Fig. 1, where we ease the notation and use the short-hand $\sigma_{c}=\sigma$ for the colloid diameter, $Q_{c}=Z$ for the colloid charge number, and $\sigma_{+}=\sigma_{-}=\sigma_{i}$ for the ion diameter that is assumed to be equal for the cations and the anions.
Figure 1: (a) Primitive model of a colloidal dispersion in a 1:1 electrolyte, represented as a three-component mixture of charged colloids, counterions, and coions with charges $Zq$, $-q$, and $+q$, and diameters $\sigma$, $\sigma_{i}$, and $\sigma_{i}$, respectively. (b) Typical configuration from a primitive model simulation with $N_{c}=32$ colloidal particles of valency $Z=90$, $N_{+}=2551$ coions, $N_{-}=5431$ counterions, and size ratio $\sigma_{i}/\sigma=0.05$. This snapshot is representative of the simulations used to generate the training data.
@@
$U_{WCA}(r_{kl})=\begin{cases}4\overline{\varepsilon}\left[\left(\frac{\sigma_{kl}}{r_{kl}}\right)^{12}-\left(\frac{\sigma_{kl}}{r_{kl}}\right)^{6}\right]+\overline{\varepsilon},&\ r_{kl}\leq 2^{1/6}\sigma_{kl};\\ 0,&\ r_{kl}>2^{1/6}\sigma_{kl},\end{cases}$ (1)
@@
The thermodynamic equilibrium properties of this primitive model are governed by six dimensionless parameters, which reduce to five by considering equal ion diameters $\sigma_{+}=\sigma_{-}\equiv\sigma_{i}$. Using the shorthand notation $N_{c}=N$, the first four key parameters are (i) the colloid packing fraction $\eta=(\pi\sigma^{3}/6)N/V$, (ii) the ion-to-colloid size ratio $\sigma_{i}/\sigma$, (iii) the colloid-to-ion charge asymmetry or colloid valency $Z$, and (iv) the size ratio $\sigma/\lambda_{B}$, which plays the role of an effective (dimensionless) temperature. In addition, we require a fifth dimensionless parameter to characterize the salt concentration, which can be expressed in various ways, such as the ion packing fraction $(\pi\sigma_{i}^{3}/6)(N_{+}+N_{-})/V$, the ion activity $\exp(\beta\mu)$, or the ion number ratio $N_{+}/N_{-}$, which varies from 0 in the salt-free case to unity in salt-dominated systems. In this work, we opt for the $\beta\mu$-dependent parameter (v) $\kappa\sigma$, where $\kappa$ is the inverse Debye screening length of the colloid-free salt reservoir in osmotic equilibrium with the colloidal suspension. This implies that the chemical potential of the salt in the reservoir is equal to that in the suspension. At a fixed salt (ion) chemical potential $\mu$, or more precisely $\beta\mu$, the inverse Debye screening length in the salt reservoir is given by $\kappa=\sqrt{4\pi\lambda_{B}c_{s}}$, where $c_{s}=(n_{+}+n_{-})$ is the salt concentration in the salt reservoir, consisting of counterion concentration $n_{+}$ and coion concentration $n_{-}$. In the reservoir, these concentrations are equal, so $n_{+}=n_{-}$. It is important to note that this reservoir concentration differs from the total (thermally averaged) salt concentration in the suspension $\big{(}N_{+}(\mu,\eta)+N_{-}(\mu,\eta)\big{)}/V$, or the added salt concentration $2N_{+}(\mu,\eta)/V$. However, these quantities coincide in the colloid-dilute limit $\eta\rightarrow 0$ or in the high-salinity limit $N_{+}/N_{-}\rightarrow 1$.
@@ II.2 Training Data Generation
The training data required to construct an effective machine-learned (ML) colloidal potential is obtained from molecular dynamics (MD) simulations using the primitive model. In particular, the ML potential is trained on the average PM force ${\bf F}^{PM}_{i}$ on colloid $i=1,\dots,N$ in specific colloid configurations $\{{\bf R}\}$, where the averaging is performed over the ion degrees of freedom of the primitive model. Each training set consists of configurations sampled at colloid packing fractions within the range $\eta\in[0.001,0.45]$, while keeping $Z$, $\sigma_{i}/\sigma$, $\sigma/\lambda_{B}$, and $\beta\mu$ fixed. To initialize a simulation with given particle sizes, colloid valency, and temperature, we begin by randomly placing $N$ colloids and $N_{-}=ZN$ charge-neutralizing counterions in a cubic box of length $L=V^{1/3}$, while preventing hard-sphere overlaps between particles and where $V$ is chosen such that the desired packing fraction $\eta=(\pi\sigma^{3}/6)N/L^{3}$ is obtained. The data-generating primitive model simulations involve a fixed number of $N$ colloids and are performed using the LAMMPS software package.Thompson et al. (2022) The equations of motion are integrated using Verlet integration, and the Nosé-Hoover thermostat ensures that the temperature stays constant at $k_{B}T/\epsilon=1.0$.Frenkel and Smit (2002) The long-range Coulombic interactions are computed using the particle-particle-particle-mesh (PPPM) Ewald summation method.Hockney and Eastwood (1988) Periodic boundary conditions are applied to mimic bulk behavior. Below, we distinguish between the salt-free limit $\exp{[\beta\mu]}\rightarrow 0$, which we will also refer to as the asymmetric electrolyte because only colloids and counterions are present, and the case with added salt, where pairs of counterions and coions are inserted into or removed from the simulation box using a Grand Canonical Monte Carlo scheme controlled by the salt chemical potential $\beta\mu$.Frenkel and Smit (2002)
@@
[…]
$\frac{V}{(N_{s}+2)\Lambda^{3}}$ $\exp{\big{[}\beta(\mu-\mathcal{U}(N_{s}+2)+\mathcal{U}(N_{s}))\big{]}}\Big{]};$
[…]
$\frac{N_{s}\Lambda^{3}}{V}\exp{\big{[}-\beta(\mu+\mathcal{U}(N_{s}-2)-\mathcal{U}(N_{s}))\big{]}}\Big{]},$
@@ II.3 Machine Learning Procedure
We proceed with an ML procedure to construct a ML potential energy function $U^{ML}(\{\bf{R}\})$ that depends on the colloid coordinates and represents a free energy (or grand potential). The scheme is designed such that the ML force $-\nabla_{i}U^{ML}(\{{\bf R}\})$ accurately matches the ion-averaged PM force ${\bf F}^{PM}_{i}$ acting on each colloid $i=1,\dots,N$ in configuration $\{{\bf R}\}$, as obtained from primitive model simulations, where the operator $\nabla_{i}$ denotes the gradient with respect to the center-of-mass position ${\bf R}_{i}$ of colloid $i$. Throughout the ML procedure, we assume that the potential energy function can be expressed as a sum of individual colloid contributions $U^{ML}(\{{\bf R}\})=\sum_{i=1}^{N}U^{ML}_{i}(\{{\bf R}\})$, where each energy contribution $U^{ML}_{i}$ is expressed in terms of a set of descriptors that characterize the local (rather than global) environment of colloid $i$ in configuration $\{{\bf R}\}$. In this work, we employ the atom-centered symmetry functions introduced by Behler and Parrinello Behler and Parrinello (2007) to represent the local environment of each colloid, where the locality is characterised by a maximum cut-off distance $R_{c}$. We distinguish between two-body and three-body symmetry functions: the two-body terms characterize the environment of particle $i$ solely through distances $R_{ij}=|{\bf R}_{j}-{\bf R}_{i}|$ to nearby colloids $j$, while the three-body terms additionally incorporate angular dependencies. Specifically, to characterize the local environment of colloid $i$ within a cutoff distance $R_{c}$ at the two-body level, we use symmetry functions
@@
where the sum runs over all distinct pairs $j,k\neq i$ within the cut-off distance $R_{c}$ to particle $i$ and $\theta_{ijk}$ is the angle between the vectors ${\bf R}_{ij}={\bf R}_{j}-{\bf R}_{i}$ and ${\bf R}_{ik}={\bf R}_{k}-{\bf R}_{i}$. The optimisation parameters $\gamma$, and $\lambda$ control the radial and angular resolution, respectively, and the parameter $\xi$ can only take values of $+1$ or $-1$.
@@
where $j=1,...,N$ labels the colloidal particles at positions $\{\mathbf{R}\}$ and $D$ is the total number of employed symmetry functions. As mentioned before, the effective ML force on particle $i$ is then obtained by taking the (negative) gradient of $U^{ML}$ with respect to position $\mathbf{R}_{i}$, which yields
@@
where $\bar{\mathbf{f}}^{PM}$ is a vector of dimension $3NM$ in which each entry is equal to the mean of all components of $\mathbf{f}^{PM}$, and thus $\bar{\mathbf{f}}^{PM}\cdot\mathbf{1}=\mathbf{f}^{PM}\cdot\mathbf{1}$, where $\mathbf{1}$ denotes the $3NM$-dimensional vector with all entries equal to one. The coefficient of determination quantifies how well the ML forces reproduce the input forces from simulations, with $R^{2}=1$ indicating perfect agreement, corresponding to an RMSE of 0.

### v2

@@ Abstract
Effective interactions between charged particles dispersed in an electrolyte are most commonly modeled using the Derjaguin-Landau-Verwey-Overbeek (DLVO) potential, where the ions in the suspension are coarse-grained out at mean-field level. However, several experiments point to shortcomings of this theory, as the distribution of ions surrounding colloids is governed by nontrivial correlations in regimes of strong Coulomb coupling (e.g. low temperature, low dielectric constant, high ion valency, high surface charge). Insight can be gained by explicitly including the ions in simulations of these colloidal suspensions, even though direct simulations of dispersions of highly charged spheres are computationally demanding. To circumvent slow equilibration, we employ a machine-learning (ML) framework to generate density-dependent ML potentials that accurately describe the effective colloid interactions at given system parameters. These ML potentials enable fast simulations and make large-scale simulations of charged colloids in suspension possible, opening the possibility for a systematic study of their phase behaviour, in particular gas-liquid and fluid-solid coexistence.
@@ I Introduction
When the refractive index of the solvent is matched to that of the colloidal particles, the attractive part of the DLVO potential, arising from Van der Waals forces due to dipole-dipole fluctuations, can be neglected. Under these conditions, the interaction between two colloidal particles with the same charge sign reduces to a purely repulsive electrostatic potential. It was therefore surprising when a series of experiments reported attractive interactions between like-charged colloids in systems of highly charged, micron-scale particles at low-salt concentrations,Kepler and Fraden (1994); Ito, Yoshida, and Ise (1994); Tata et al. (1997); Larsen and Grier (1997); Gomez et al. (2009); Wang et al. (2024) challenging the universal validity of DLVO theory. It should be noted, however, that the results of some of these experiments remain controversial, due to issues related to reproducibility Palberg and Würth (1994) and potential experimental artifacts.Squires and Brenner (2000) This low-salt regime is characterized by long screening lengths, which increases the importance of many-body interactions beyond simple pairwise interactions, as the larger electric double layers more frequently overlap with multiple neighboring EDLs. A second characteristic of this regime is that, as the colloid concentration increases, the ionic strength, and thus the effective Debye length, is no longer determined predominantly by the concentration of the background electrolyte, but rather by the higher concentration of counterions. Various studies based on Poisson-Boltzmann calculations have proposed that these two effects are the leading causes behind like-charge attractions—specifically, gas-liquid and gas-crystal phase separation—in low-salt colloidal suspensions.van Roij and Hansen (1997); van Roij, Dijkstra, and Hansen (1999); Zoetekouw and van Roij (2006); Warren (2000); Russ et al. (2002) A second cause of like-charged attractive interactions lies in the strongly coupled electrostatic regime, which is characterized by low temperature, low dielectric constant, and high ionic valency.Kjellander, Marčelja, and Quirk (1988); Allahyarov, D’Amico, and Löwen (1998); Naji et al. (2005); Punkkinen et al. (2008) A defining feature of this regime is the binding of point-like ions onto highly charged surfaces, forming a quasi-two-dimensional layer rather than the diffuse three-dimensional distribution characteristic of the Poisson-Boltzmann electric double layer. When two such decorated surfaces approach each other, strong ion-ion correlations give rise to attractive interactions.
@@
In this paper, we also coarse-grain the ions to significantly reduce simulation times, but we achieve this by employing a machine learning (ML) framework to generate effective many-body colloidal potentials. While ML approaches have been successfully used to accelerate simulations of uncharged systems,Deringer, Caro, and Csányi (2019); Campos-Villalobos et al. (2021); Nguyen and Huang (2022); Argun, Fu, and Statt (2024) the present work is the first to apply machine-learned effective potentials to charged colloids. To construct these ML colloidal potentials, we adopt a linear regression scheme previously used to develop effective potentials for sterically stabilized colloids with non-adsorbing polymers Campos-Villalobos et al. (2021) and for ligand-stabilized colloids.Giunta, Campos-Villalobos, and Dijkstra (2023) Here, we train the ML potentials using data from primitive model simulations, resulting in effective potentials that accurately capture the effective interactions between charged colloids in the presence of ions, in addition to bare-colloid interactions. In subsequent “colloids-only” simulations using these ML potentials, the ions are effectively integrated out. Moreover, in this representation, the simulations do not include any explicit charges, since the colloids interact through the effective coarse-grained ML potential, which eliminates the need to account for long-range interactions. These two factors bring about a significant increase in computational efficiency –at least twenty times faster for colloidal systems–, enabling studies at higher colloid charges and salt concentrations.
@@ II.1 Primitive Model simulations
The system of interest in this study is a three-dimensional primitive model (PM) of a colloidal dispersion in a 1:1 electrolyte. This colloidal suspension is viewed as a three-component mixture of charged colloids, monovalent counterions, and monovalent coions, while the solvent is modeled as a structureless dielectric continuum. The colloidal particles have an effective hard-core diameter $\sigma_{c}$ and a total positive charge $Q_{c}q$ uniformly distributed over their surfaces, with valency $Q_{c}\geq 1$ and elementary charge $q$. The monovalent counterions are anions with diameter $\sigma_{-}$ and charge q-q, and the monovalent coions are cations of diameter $\sigma_{+}$ and charge +q+q. These three species are illustrated in Fig. 1, where we ease the notation and use the short-hand $\sigma_{c}=\sigma$ for the colloid diameter, $Q_{c}=Z$ for the colloid charge number, and $\sigma_{+}=\sigma_{-}=\sigma_{i}$ for the ion diameter that is assumed to be equal for the cations and the anions.
Figure 1: (a) Primitive model of a colloidal dispersion in a 1:1 electrolyte, represented as a three-component mixture of charged colloids, counterions, and coions with charges ZqZq, q-q, and +q+q, and diameters $\sigma$, $\sigma_{i}$, and $\sigma_{i}$, respectively. (b) Typical configuration from a primitive model simulation with $N_{c}=32$ colloidal particles of valency $Z=90$, $N_{+}=2551$ coions, $N_{-}=5431$ counterions, and size ratio $\sigma_{i}/\sigma=0.05$. This snapshot is representative of the simulations used to generate the training data.
@@
$\hskip-5.69054ptU_{WCA}(r_{kl})=\begin{cases}4\overline{\varepsilon}\left[\left(\frac{\sigma_{kl}}{r_{kl}}\right)^{12}\hskip-5.69054pt-\left(\frac{\sigma_{kl}}{r_{kl}}\right)^{6}\right]+\overline{\varepsilon},&\ \hskip-5.69054ptr_{kl}\leq 2^{1/6}\sigma_{kl};\\ 0,&\ \hskip-5.69054ptr_{kl}>2^{1/6}\sigma_{kl},\end{cases}$ (1)
@@
The thermodynamic equilibrium properties of this primitive model are governed by six dimensionless parameters, which reduce to five by considering equal ion diameters $\sigma_{+}=\sigma_{-}\equiv\sigma_{i}$. Using the shorthand notation $N_{c}=N$, the first four key parameters are (i) the colloid packing fraction $\eta=(\pi\sigma^{3}/6)N/V$, (ii) the ion-to-colloid size ratio $\sigma_{i}/\sigma$, (iii) the colloid-to-ion charge asymmetry or colloid valency $Z$, and (iv) the size ratio $\sigma/\lambda_{B}$, which plays the role of an effective (dimensionless) temperature. In addition, we require a fifth dimensionless parameter to characterize the salt concentration, which can be expressed in various ways, such as the ion packing fraction $(\pi\sigma_{i}^{3}/6)(N_{+}+N_{-})/V$, the ion activity $\exp(\beta\mu)$, or the ion number ratio $N_{+}/N_{-}$, which varies from 0 in the salt-free case to unity in salt-dominated systems. In this work, we opt for the $\beta\mu$-dependent combination $\kappa\sigma$ as the fifth dimensionless parameter, where $\kappa$ is the inverse Debye screening length of the colloid-free salt reservoir in osmotic equilibrium with the colloidal suspension. This implies that the chemical potential of the salt in the reservoir is equal to that in the suspension. At a fixed salt (ion) chemical potential $\mu$, or more precisely $\beta\mu$, the inverse Debye screening length in the salt reservoir is given by $\kappa=\sqrt{4\pi\lambda_{B}c_{s}}$, where $c_{s}=(n_{+}+n_{-})$ is the salt concentration in the salt reservoir, consisting of counterion concentration $n_{+}$ and coion concentration $n_{-}$. In the reservoir, these concentrations are equal, so $n_{+}=n_{-}$. It is important to note that this reservoir concentration differs from the total (thermally averaged) salt concentration in the suspension $\big(N_{+}(\mu,\eta)+N_{-}(\mu,\eta)\big)/V$, or the added salt concentration $2N_{+}(\mu,\eta)/V$. However, these quantities coincide in the colloid-dilute limit $\eta\rightarrow 0$ or in the high-salinity limit $N_{+}/N_{-}\rightarrow 1$.
@@ II.2 Training Data Generation
The training data required to construct an effective machine-learned (ML) colloidal potential is obtained from molecular dynamics (MD) simulations using the primitive model. In particular, the ML potential is trained on the average PM force ${\bf F}^{PM}_{i}$ on colloid $i=1,\dots,N$ in specific colloid configurations $\{{\bf R}\}$, where the averaging is performed over the ion degrees of freedom of the primitive model. Each training set consists of configurations sampled at colloid packing fractions within the range $\eta\in[0.001,0.45]$, while keeping $Z$, $\sigma_{i}/\sigma$, $\sigma/\lambda_{B}$, and $\beta\mu$ fixed. Our goal is to generate an ML potential that can be used to simulate a system characterized by $Z$, $\sigma_{i}/\sigma$, $\sigma/\lambda_{B}$, and $\beta\mu$ across a range of packing fractions, thereby enabling the study of not only single bulk phases at varying densities but also phase-separated systems with different densities. It should be noted, however, that a separate ML potential must be trained for each distinct set of system parameters ($Z$, $\sigma_{i}/\sigma$, $\sigma/\lambda_{B}$, $\beta\mu$).
To initialize a simulation with given particle sizes, colloid valency, and temperature, we begin by randomly placing $N$ colloids and $N_{-}=ZN$ charge-neutralizing counterions in a cubic box of length $L=V^{1/3}$, while preventing hard-sphere overlaps between particles and where $V$ is chosen such that the desired packing fraction $\eta=(\pi\sigma^{3}/6)N/L^{3}$ is obtained. The data-generating primitive model simulations involve a fixed number of $N$ colloids and are performed using the LAMMPS software package.Thompson et al. (2022) The equations of motion are integrated using the Verlet algorithm, and the Nosé-Hoover thermostat is employed to maintain a constant temperature of $k_{B}T/\epsilon=1.0$, with a relaxation time of $0.05\tau_{MD}$, where $\tau_{MD}=\sqrt{m\sigma^{2}/(k_{B}T)}$.Frenkel and Smit (2002) The long-range Coulombic interactions for distances $R>2.5\sigma$ are computed using the particle-particle-particle-mesh (PPPM) Ewald summation method, with a target relative force error of $10^{-4}k_{B}T/\sigma$.Hockney and Eastwood (1988) Periodic boundary conditions are applied to mimic bulk behavior. Below, we distinguish between the salt-free limit $\exp{[\beta\mu]}\rightarrow 0$, which we will also refer to as the asymmetric electrolyte because only colloids and counterions are present, and the case with added salt, where pairs of counterions and coions are inserted into or removed from the simulation box using a Grand Canonical Monte Carlo scheme controlled by the salt chemical potential $\beta\mu$.Frenkel and Smit (2002)
@@
$\text{acc}(N_{s}\rightarrow N_{s}+2)=$ $\text{min}\Big[1,$ (3)
$\hskip-14.22636pt\frac{V}{(N_{s}+2)\Lambda^{3}}$ $\exp{\big[\beta(\mu-\mathcal{U}(N_{s}+2)+\mathcal{U}(N_{s}))\big]}\Big];$
$\text{acc}(N_{s}\rightarrow N_{s}-2)=$ $\text{min}\Big[1,$ (4)
$\hskip 0.0pt\hskip-42.67912pt\frac{N_{s}\Lambda^{3}}{V}\exp{\big[-\beta(\mu+\mathcal{U}(N_{s}-2)-\mathcal{U}(N_{s}))\big]}\Big],$
@@ II.3 Machine Learning Procedure
We proceed with an ML procedure to construct a ML potential energy function $U^{ML}(\{\bf{R}\})$ that depends on the colloid coordinates and represents a free energy (or grand potential). The scheme is designed such that the ML force $\nabla_{i}U^{ML}(\{{\bf R}\})$ accurately matches the ion-averaged PM force ${\bf F}^{PM}_{i}$ acting on each colloid $i=1,\dots,N$ in configuration $\{{\bf R}\}$, as obtained from primitive model simulations, where the operator $\nabla_{i}$ denotes the gradient with respect to the center-of-mass position ${\bf R}_{i}$ of colloid $i$. Throughout the ML procedure, we assume that the potential energy function can be expressed as a sum of individual colloid contributions $U^{ML}(\{{\bf R}\})=\sum_{i=1}^{N}U^{ML}_{i}(\{{\bf R}\})$, where each energy contribution $U^{ML}_{i}$ is expressed in terms of a set of descriptors that characterize the local (rather than global) environment of colloid $i$ in configuration $\{{\bf R}\}$. In this work, we employ the atom-centered symmetry functions introduced by Behler and Parrinello Behler and Parrinello (2007) to represent the local environment of each colloid, where the locality is characterised by a maximum cut-off distance $R_{c}$. We distinguish between two-body and three-body symmetry functions: the two-body terms characterize the environment of particle $i$ solely through distances $R_{ij}=|{\bf R}_{j}-{\bf R}_{i}|$ to nearby colloids $j$, while the three-body terms additionally incorporate angular dependencies. Specifically, to characterize the local environment of colloid $i$ within a cutoff distance $R_{c}$ at the two-body level, we use symmetry functions
@@
where the sum runs over all distinct pairs $j,k\neq i$ within the cut-off distance $R_{c}$ to particle $i$ and $\theta_{ijk}$ is the angle between the vectors ${\bf R}_{ij}={\bf R}_{j}-{\bf R}_{i}$ and ${\bf R}_{ik}={\bf R}_{k}-{\bf R}_{i}$. The optimisation parameters $\gamma$, and $\lambda$ control the radial and angular resolution, respectively, and the parameter $\xi$ can only take values of +1+1 or 1-1.
@@
where j=1,,Nj=1,...,N labels the colloidal particles at positions $\{\mathbf{R}\}$ and $D$ is the total number of employed symmetry functions. As mentioned before, the effective ML force on particle $i$ is then obtained by taking the (negative) gradient of $U^{ML}$ with respect to position $\mathbf{R}_{i}$, which yields
@@
where $\bar{\mathbf{f}}^{PM}$ is a vector of dimension $3NM$ in which each entry is equal to the mean of all components of $\mathbf{f}^{PM}$, and thus $\bar{\mathbf{f}}^{PM}\cdot\mathbf{1}=\mathbf{f}^{PM}\cdot\mathbf{1}$, where $\mathbf{1}$ denotes the $3NM$-dimensional vector with all entries equal to one. The coefficient of determination quantifies how well the ML forces reproduce the input forces from simulations, with $R^{2}=1$ indicating perfect agreement, corresponding to an RMSE of 0. Once $R^{2}$ approaches $1$ and the forces predicted by the ML potential closely match those from the training data across their full range, we proceed by performing simulations using the ML potential.